# RAIL: Robot Affordance Imagination with Large Language Models

### v1

@@ III-D Affordance Evaluation
In this part, the algorithm analyzes the results of the simulation $\mathcal{R}$ to evaluate the object affordance. It consists of two LLM modules to evaluate $r$ of each plan and the object affordance, respectively.
@@ III-D2 Functional Pose Analyzer
BLIP functional pose 45.2 13.0 95.2 70.6 11.1 85.1 87.2 68.2 56.8
@@ IV Experiments
We implement our framework system with Python, using Pybullet [28] as the physics simulator. The algorithm is evaluated on a computer running an Intel Core i7-11370H @ 3.3GHz CPU and Nvidia GTX 3060 GPU. In a real experiment setting Fig 5 (b), a Franka Emika robot arm is used for manipulation. An RGBD camera is mounted on the end effector for reconstruction. The language model in this study is founded on GPT-4 [29]. To enhance response time and speed of generation, we opt for the GPT-4-turbo variant and set the parameter Temperature = 0.1 to ensure optimality while providing a certain degree of generalization.
@@ IV-A Data
Our dataset comprises 301 synthetic and 19 real-life objects. The synthetic objects, sourced from a subset of the Princeton ModelNet40 dataset [30], span 8 classes: cup, basket, bathtub, chair, plate, table, vase and bowl. To optimize the performance of LLM modules and the simulator, we utilize 6 synthetic objects across 3 classes, i.e., cup, vase, and chair. The rest data, which are unseen by our system, are used as the test set. To assess affordance classification, non-functional objects from various classes are also included in the testing phase.
In real robot experiments, we use a cup to tune the parameters for the robot planning to shorten the sim-to-real gap. The remaining 18 novel objects, varied in size and appearance, are used to test the recognization of 15 classes of affordances and affordance-based manipulation. This includes 13 novel affordances previously unencountered by our system.
@@ IV-B Real Robot Experiment
The robot arm first moves to 12 predefined poses to scan the object. The object model is cropped and segmented from the scene point cloud reconstructed using TSDF-Fusion [31]. With the requested affordance name segmented and passed to the affordance reasoning module, the algorithm imagines the object model and classifies if the object can be used with the requested affordance in the placed pose. If the object is recognized as functional and in a functional pose, the algorithm proposes an optimal agent trajectory. The robot then requests a volunteer to hand over the real agent. Holding the real agent, the robot executes the imagined trajectory and positions the agent, transferring the affordance understanding to practical actions. Throughout this process, simple force control is used for robot manipulation, pausing its motion if external force surpasses a set limit, ensuring safety and precision in task execution.
@@ IV-C Baseline
We compare our method with the baseline, BLIP [32], which is a Visiual Question Answering (VQA) model that makes responses to user’s prompt, we make comparisons on two tasks: affordance classification and functional pose prediction.
For affordance classification, we present an object’s image alongside the query “Is it a (requested affordance name)?”, and it ascertains whether the model possesses the requested affordance. To study how the object’s pose impacts its performance and demonstrate the versatility of our approach across different poses, for each data sample we capture two images in arbitrary and functional poses, respectively. In functional pose prediction, we change the query to ”Can the (requested affordance name) function in this pose?” With the image input, it makes judgement whether the object is in its functional pose. By doing so, we evaluate its ability to make responses on object pose information by visual cues.
@@ V-A Affordance Classification
The affordance classification achieves high success rates on synthetic data, as shown in Tab. I. Notably, the classification of novel affordance classes demonstrates exceptional performance, achieving a 88.2% success rate. Failure examples are mainly due to improperly generated profiles, such as imagination outline analysis, agent’s model, trajectory parameters of agent motion, etc. We notice that the performance is degraded on vases, which we hypothesize is because the geometry varies greatly and the openings are usually small. In the absence of specific information about geometry, the generated agent may only match the overall dimensions of the object and it might be oversized for the opening, leading to a failure of the affordance imagination. BLIP does not perform as well on synthetic data even with objects in a functional pose. It shows good performances only for classes with unique appearance features such as chairs and vases. When objects are placed in random poses, it results in an overall correctness of only 30.2%.
@@ V-B Functional Pose Prediction
We evaluate the functional pose prediction results on objects that are successfully recognized as functional. In Tab II, our method achieves very high accuracy in functional pose prediction for synthetic data. When considering only the current pose, the robot successfully recognizes the object as being in a functional pose across all real-world trials.
@@ V-C Real Robot Manipulation
Qualitative results are shown in Fig. 7, our system achieves 100% across 20 novel trials encompassing 15 distinct tasks, utilizing 18 objects previously unknown to the system. Notably, the system exhibits impressive generalization capabilities by successfully recognizing and manipulating 13 new affordances. Furthermore, it can identify objects with multifunctional uses and adapts its interaction accordingly. For example, the rightmost two trials in Fig. 7 showcase an object that the system recognizes both as a shelf, by positioning the agent on top, and as a cupboard, by inserting the agent inside.

### v2

@@ III-D Affordance Evaluation
In this part, the algorithm analyzes the results of the simulation $\mathcal{R}$ to evaluate the object affordance. It consists of two LLM modules to evaluate $r$ of each plan and the overall object affordance, respectively.
@@ III-D2 Functional Pose Analyzer
BLIP [35] functional pose 45.2 13.0 95.2 70.6 11.1 85.1 87.2 68.2 56.8
@@ IV Experiments
We implement our framework system with Python, using Pybullet [36] as the physics simulator. The algorithm is evaluated on a computer running an Intel Core i7-11370H @ 3.3GHz CPU and Nvidia GTX 3060 GPU. In a real experiment setting Fig. 5 (b), a Franka Emika robot arm is used for manipulation. An RGBD camera is mounted on the end effector for reconstruction. The language model in this study is founded on GPT-4 [37]. To enhance response time and speed of generation, we opt for the GPT-4-turbo variant and set the parameter Temperature = 0.1 to ensure optimality while providing a certain degree of generalization.
@@ IV-A Data
Our dataset comprises 301 synthetic and 19 real-life objects. The synthetic objects, sourced from a subset of the Princeton ModelNet40 dataset [38], span 8 classes: cup, basket, bathtub, chair, plate, table, vase, and bowl. To optimize the performance of LLM modules and the simulator, we utilize 6 synthetic objects across 3 classes, i.e., cup, vase, and chair, to tune the prompt. The rest data, which are unseen by our system, are used as the test set. To assess affordance classification, non-functional objects from various classes are also included in the testing phase.
In real robot experiments, we use a cup to tune the parameters for the robot planning to shorten the sim-to-real gap. The remaining 18 novel objects, varied in size and appearance, are used to test the recognition of 15 classes of affordances and affordance-based manipulation. This includes 13 novel affordances previously unencountered by our system.
@@ IV-B Real Robot Experiment
The robot arm first moves to 12 predefined poses to scan the object. The object model is cropped and segmented from the scene point cloud reconstructed using TSDF-Fusion [39]. With the requested affordance name segmented and passed to the affordance reasoning module, the algorithm imagines the object model and classifies if the object can be used with the requested affordance in the placed pose. If the object is recognized as functional and in a functional pose, the algorithm proposes an optimal agent trajectory. The robot then requests a volunteer to hand over the real agent. Having the real agent in hand, the robot executes the imagined trajectory and positions the agent, transferring the affordance understanding to practical actions. Throughout this process, simple force control is used for robot manipulation, pausing its motion if external force surpasses a pre-defined limit, ensuring safety and precision in task execution.
@@ IV-C Baseline
We compare our method with the baseline, BLIP [35], which is a unified vision-language model for image understanding and generation, and we use the version fined-tuned on Visual Question Answering (VQA) task that makes responses to user’s prompt, we make comparisons on two criteria: affordance classification and functional pose prediction.
For affordance classification, we present an object’s image alongside the query “Is it a (requested affordance name)?”, and it ascertains whether the model possesses the requested affordance. In functional pose prediction, we input the image together with the query ”Can the (requested affordance name) function in this pose?” To study how the object’s pose impacts its performance and demonstrate the versatility of our approach across different poses, for each data sample we capture two images in arbitrary and functional poses, respectively.
@@ V-A Affordance Classification
The affordance classification achieves high success rates on synthetic data, as shown in Tab. I. Notably, the classification of novel affordance classes demonstrates exceptional performance, achieving an 88.2% success rate. Failure examples are mainly due to improperly generated profiles, such as imagination outline analysis, agent’s model, trajectory parameters of agent motion, etc. We notice that the performance is degraded on vases, which we hypothesize is because the geometry varies greatly and the openings are usually small. In the absence of specific information about geometry, the generated agent may only match the overall dimensions of the object and it might be oversized for the opening, leading to a failure of the affordance imagination. BLIP does not perform as well on synthetic data even with objects in a functional pose. It shows good performances only for classes with unique appearance features such as chairs and vases. When objects are placed in random poses, it results in an overall correctness of only 30.2%.
@@ V-B Functional Pose Prediction
We evaluate the functional pose prediction results on objects that are successfully recognized as functional. In Tab. II, our method achieves very high accuracy in functional pose prediction for synthetic data. When considering only the current pose, the robot successfully recognizes the object as being in a functional pose across all real-world trials.
@@ V-C Real Robot Manipulation
Qualitative results are shown in Fig. 7, our system achieves 100% across 20 novel trials encompassing 15 distinct tasks, utilizing 18 objects previously unknown to the system. Notably, the system exhibits impressive generalization capabilities by successfully recognizing and manipulating 13 new affordances. Furthermore, it can identify objects with multi-functional usage and adapts its interaction accordingly. For example, the rightmost two trials in Fig. 7 showcase an object that is recognized as possessing two types of affordances, i.e., a shelf by positioning the agent on top, and a cupboard by inserting the agent inside.